# Bijective rigidity of uniform Roe algebras and injectivity of the comparison map

### v1

@@ Introduction
Coarse geometry studies metric spaces by focusing on their large-scale structure while disregarding small-scale features. Since local properties play no essential role in this setting, it is customary to restrict attention to discrete metric spaces. Throughout this paper, we work with uniformly locally finite metric spaces (see Definition 1.1), also referred to as bounded geometry metric spaces. A classic example is provided by a Cayley graph of a finitely generated group with the shortest path metric. Let $(X,d)$ and $(Y,\partial)$ be uniformly locally finite metric spaces. A map $f\colon X\to Y$ is called coarse if for every $R\geq 0$ there exists $S\geq 0$ such that $f$ maps subsets of $X$ of diameter at most $R$ to subsets of $Y$ of diameter at most $S$. Coarse maps are precisely those maps that preserve the large-scale geometry of metric spaces. Two coarse maps $f,g\colon X\to Y$ are said to be close if they are at bounded distance, that is,
@@
The coarse category $\mathbf{Coarse}$ is the category whose objects are uniformly locally finite metric spaces and whose morphisms are closeness classes of uniformly finite-to-one [^1] A map $f\colon X\to Y$ is uniformly finite-to-one if there exists $N\geq 1$ such that the preimage of every point of $Y$ consists of at most $N$ elements (see Section 1.1). Coarse equivalences are automatically uniformly finite-to-one. coarse maps. Isomorphisms in $\mathbf{Coarse}$ are called coarse equivalences. In the case of finitely generated groups equipped with word metrics, the notion of coarse equivalence coincides with quasi-isometry. We say that two uniformly locally finite metric spaces $(X,d)$ and $(Y,\partial)$ are (bijectively) coarsely equivalent if there exists a (bijective) coarse equivalence between them.
Operator-algebraic methods enter coarse geometry through the association of a $C^{*}$-algebra $C^{*}_{u}(X)$, called the uniform Roe algebra, to a uniformly locally finite metric space $(X,d)$. These $C^{*}$-algebras are variants of the Roe algebra introduced by Roe in [19] in his study of index theory on non-compact Riemannian manifolds. More precisely, let $(X,d)$ be a uniformly locally finite metric space. A bounded operator $T\in\mathcal{B}(\ell^{2}(X))$ is said to have controlled propagation if the quantity
@@
is finite. The uniform Roe algebra $C^{*}_{u}(X)$ is defined as the norm closure in $\mathcal{B}(\ell^{2}(X))$ of the *-algebra of bounded operators with controlled propagation. For instance, if $G$ is a finitely generated discrete group equipped with a word metric, then the uniform Roe algebra of $G$ is canonically isomorphic to the reduced crossed product $\ell^{\infty}(G)\rtimes_{r}G$. Uniform Roe algebras encode a substantial amount of coarse-geometric information about the underlying space. For example, a uniformly locally finite metric space $(X,d)$ has property A if and only if $C^{*}_{u}(X)$ is a nuclear $C^{*}$-algebra [23], and $(X,d)$ is amenable (as a metric space) if and only if $C^{*}_{u}(X)$ admits a tracial state [1]. This naturally leads to the question of whether the uniform Roe algebra is a complete invariant of the coarse geometry of the space.
Conjecture 0.1.
@@
Let $(X,d)$ and $(Y,\partial)$ be uniformly locally finite metric spaces. Then the $C^{*}$-algebras $C^{*}_{u}(X)$ and $C^{*}_{u}(Y)$ are Morita equivalent (respectively, isomorphic) if and only if $(X,d)$ and $(Y,\partial)$ are coarsely equivalent (respectively, bijectively coarsely equivalent).
Conjecture 0.1 is commonly referred to as the (bijective) rigidity problem. In [24], it was shown that uniformly locally finite metric spaces with property A satisfy the rigidity problem. A positive answer to the bijective rigidity problem for property A spaces was subsequently obtained in [27]. A complete solution to the rigidity problem was later established in [2], and further generalised to a broader class of Roe-type algebras in [14, 15]. Taken together, these results yield the following theorem.
Theorem 0.1 ([2, Theorem 1.4]).
Expander graphs provide fundamental examples of uniformly locally finite metric spaces that do not satisfy property A. In [3], it was shown that the bijective rigidity problem holds for expander graphs. One of the main techniques involved is uniformly finite homology (see Definition 1.10), a homology theory for uniformly locally finite metric spaces introduced by Block and Weinberger [5]. The degree-zero uniformly finite homology group $H_{0}^{\mathrm{uf}}(X;\mathbb{Z})$ detects whether a coarse equivalence is close to a bijective coarse equivalence, as made precise by the following theorem.
Theorem 0.2 ([28, Theorem A]).
@@
Let $(X,d)$ and $(Y,\partial)$ be uniformly locally finite metric spaces, and let $f\colon X\to Y$ be a coarse equivalence. The following conditions are equivalent:
  1. (1)
  2. (2)
    $H^{\mathrm{uf}}_{0}(f)\bigl([\mathbbm{1}_{X}]\bigr)=[\mathbbm{1}_{Y}]$.
The theorem suggests that uniformly finite homology may play a decisive role in addressing the bijective rigidity problem. It is therefore natural to ask whether uniformly finite homology can be recovered from the uniform Roe algebra of the underlying space.
Coarse geometry is intrinsically connected to the theory of étale groupoids [^2] We refer the reader to [22] for a detailed introduction to étale groupoids and their $C^{*}$-algebras.. An étale groupoid is a topological groupoid $\mathcal{G}$ with unit space $\mathcal{G}^{(0)}$ such that the source and range maps $s,r\colon\mathcal{G}\to\mathcal{G}^{(0)}$ are local homeomorphisms. An étale groupoid $\mathcal{G}$ is said to be ample if its unit space $\mathcal{G}^{(0)}$ is totally disconnected. The theory of étale groupoids has attracted significant attention due to its deep connections with operator algebras. To any étale groupoid $\mathcal{G}$ one can associate a $C^{*}$-algebra $C^{*}_{r}(\mathcal{G})$, called the reduced groupoid $C^{*}$-algebra of $\mathcal{G}$. In [13], Li showed that every classifiable simple $C^{*}$-algebra arises as the reduced groupoid $C^{*}$-algebra of a topologically principal (possibly twisted) étale groupoid. For a uniformly locally finite metric space $(X,d)$, the authors of [23] constructed a Hausdorff, principal, ample groupoid $\mathcal{G}(X,d)$, called the coarse groupoid of $(X,d)$, whose unit space is the Stone-Cech compactification of $X$. The construction of the coarse groupoid provides a functor from the coarse category to the category of étale groupoids with generalised groupoid homomorphisms. Moreover, the reduced groupoid $C^{*}$-algebra of the coarse groupoid $\mathcal{G}(X,d)$ is canonically isomorphic to the uniform Roe algebra of $(X,d)$.
In [9], Crainic and Moerdijk introduced a homology theory for étale groupoids. In [16], Matui showed that, for groupoids arising from shifts of finite type, the $K_{0}$-groups (respectively, $K_{1}$-groups) of the associated reduced groupoid $C^{*}$-algebras are isomorphic to the direct sum of the even (respectively, odd) homology groups. He conjectured that this phenomenon extends to more general ample groupoids.
Conjecture 0.2 (HK conjecture).
$K_{*}\bigl(C^{*}_{r}(\mathcal{G})\bigr)\cong\bigoplus_{i=0}^{\infty}H_{2i+*}(\mathcal{G};\mathbb{Z}).$
The HK conjecture has been verified for many classes of ample groupoids, in some cases even beyond the stated hypotheses. However, counterexamples to Conjecture 0.2 were constructed by Scarparo [21] and Deeley [10]. It is well known that for any ample groupoid $\mathcal{G}$ and $i=0,1$ there exists canonical group homomorphism
called the $i$th comparison maps (see Definition 1.15 for the definition of $\alpha_{0}$ in the case of coarse groupoids). At present, only partial results are known concerning the existence and properties of comparison maps in higher degrees. The authors of [6] showed that the groupoid homology of a coarse groupoid is canonically isomorphic to the uniformly finite homology of the underlying uniformly locally finite metric space. Moreover, they proved that Conjecture 0.2 holds for coarse groupoids associated to uniformly locally finite metric spaces of sufficiently small asymptotic dimension.
@@
In this paper, we establish a connection between the bijective rigidity problem and the injectivity of the 0th comparison map appearing in the HK conjecture for coarse groupoids. A central obstruction is that the injectivity of the comparison map does not, in general, pass to full clopen subgroupoids. As a consequence, our approach requires working with collections of groupoids rather than with a single groupoid.
We say that a collection $\mathcal{C}$ of metric spaces is admissible (see Definition 3.1) if it is stable under taking subspaces and doublings [^3] On the level of coarse groupoids, a doubling corresponds to taking the product with $\mathcal{R}_{n}$, where $\mathcal{R}_{n}$ denotes the discrete groupoid associated to the full equivalence relation on $n$ points. See Section 1.1 for the coarse geometric definition. . For a property $\mathrm{P}$ of metric spaces, we say that an admissible collection $\mathcal{C}$ satisfies $\mathrm{P}$ if every element of $\mathcal{C}$ satisfies $\mathrm{P}$. For example, since property A passes to subspaces and is a coarse invariant, the class of metric spaces satisfying property A forms an admissible collection. We say that a class $\mathcal{C}$ of uniformly locally finite metric spaces satisfies bijective rigidity if for any uniformly locally finite metric spaces $(X,d)$ and $(Y,\partial)$ in $\mathcal{C}$ such that $C^{*}_{u}(Y)\cong C^{*}_{u}(X)$, the coarse equivalence given by Theorem 0.1 is close to a bijective coarse equivalence. The following theorem constitutes the main result of this paper.
Theorem A.
@@
    For every $(X,d)\in\mathcal{C}$, the 0th comparison map $\alpha_{0}\colon H^{\mathrm{uf}}_{0}(X;\mathbb{Z})\to K_{0}\bigl(C^{*}_{u}(X)\bigr)$ is injective;
@@
    The collection $\mathcal{C}$ satisfies bijective rigidity.
The key step in the proof of the above theorem is to show that the maps in $K$-theory induced, respectively, by an isomorphism of uniform Roe algebras and by the coarse equivalence arising from Theorem 0.1 coincide. Moreover, in Theorem A, the proof of the implication $(1)\Rightarrow(2)$ does not make use of subspaces. As a consequence, we obtain the following immediate corollary.
Corollary B.
Let $(X,d)$ and $(Y,\partial)$ be uniformly locally finite metric spaces whose uniform Roe algebras are isomorphic. If the 0th comparison map $\alpha_{0}^{X}$ is injective, then $(X,d)$ and $(Y,\partial)$ are bijectively coarsely equivalent.
Independently of the present work, Vignati proved in [26], using different methods, that the bijective rigidity problem holds for all uniformly locally finite metric spaces. Nevertheless, we formulate Theorem A and Corollary B as establishing a relationship between the two conjectural statements, as this viewpoint clarifies the techniques underlying our arguments. Moreover, we expect that the methods developed here can be adapted to broader classes of ample groupoids, potentially yielding new results on the injectivity of comparison maps and on rigidity phenomena for ample groupoids.
Since the class of all uniformly locally finite metric spaces is trivially an admissible collection, the result of [26] implies the unconditional injectivity of the 0th comparison map for coarse groupoids. On the other hand, Whyte proved in [28] that, if a uniformly locally finite metric space is coarsely connected, then its 0th uniformly finite homology group is a real vector space. Combining these two results, we obtain the following corollary.
@@
The above corollary partially extends the results of [6], where the authors proved that $\alpha_{0}$ is split injective for uniformly locally finite metric spaces of small asymptotic dimension. To the best of our knowledge, no further results concerning the injectivity of $\alpha_{0}$ for uniform Roe algebras are currently available. There exist many examples of uniformly locally finite metric spaces without property A that are not coarsely connected, such as expander graphs. By contrast, any finitely generated group equipped with a word-length metric is coarsely connected. In this setting, for a finitely generated group $G$, the uniformly finite homology of $G$ is canonically isomorphic to the group homology of $G$ with coefficients in $\ell^{\infty}(G,\mathbb{R})$.
Corollary D.
@@
Recent results of [18] show that the rational HK conjecture holds for all second-countable, locally compact, Hausdorff, ample groupoids with torsion-free stabilisers that satisfy the rational Baum–Connes conjecture. Coarse groupoids fall outside this framework: they are not second-countable and do not in general satisfy the Baum–Connes conjecture (see [12]). Nevertheless, by tensoring with $\mathbb{Q}$ in Corollary C, we immediately obtain unconditional split injectivity of the rational comparison map.
@@
In this manuscript, we do not make essential use of the theory of étale groupoids in the proofs of our results. Nevertheless, we include a discussion of the equivalent formulation in terms of coarse groupoids.
In Section 1, we introduce the necessary background from coarse geometry, uniform Roe algebras, and uniformly finite homology. We briefly recall the construction of the coarse groupoid from [23], sketch the main ideas underlying the proof of Theorem 0.1, review the results on uniformly finite homology (in particular, Theorem 0.2) of [28], and construct the 0th comparison map.
Section 2 develops a framework of covering isometries for uniform Roe algebras, analogous to the construction for Roe algebras (see [29, Section 4.3]). Let $H$ be a separable Hilbert space. A uniform cover of a uniformly finite-to-one coarse map $f\colon X\to Y$ is an isometry
@@
In Section 3, we prove Theorem A and deduce Corollary B. The key step in the proof of the implication $(1)\Rightarrow(2)$ is to show that, given an isomorphism of uniform Roe algebras $\Phi\colon C^{*}_{u}(X)\longrightarrow C^{*}_{u}(Y)$, the induced map in $K$-theory coincides with the map arising from the coarse equivalence $f_{\Phi}\colon X\longrightarrow Y$ provided by Theorem 0.1. Since $K_{0}(\Phi)$ preserves the class of the unit, it follows that
@@
modulo $\ker(\alpha_{0}^{Y})$. As the comparison maps are assumed to be injective, Theorem 0.2 completes the proof of the implication $(1)\Rightarrow(2)$. For the converse implication, to each positive function $h\in\ell^{\infty}(X,\mathbb{Z})$ we associate a subspace $X(h)$ of a doubling of $X$. We show that if $h_{1},h_{2}$ are positive functions such that
@@
This completes the proof of Theorem A. The final section discusses further consequences of Theorem A and Corollary B.

### v2

@@ Introduction
Coarse geometry studies metric spaces by focusing on their large-scale structure while disregarding small-scale features. Since local properties play no essential role in this setting, it is customary to restrict attention to discrete metric spaces. Let $(X,d)$ and $(Y,\partial)$ be metric spaces. A map $f\colon X\to Y$ is called coarse if for every $R\geq 0$ there exists $S\geq 0$ such that $f$ maps subsets of $X$ of diameter at most $R$ to subsets of $Y$ of diameter at most $S$. Coarse maps are precisely those maps that preserve the large-scale geometry of metric spaces. Two coarse maps $f,g\colon X\to Y$ are said to be close if they are at bounded distance, that is,
@@
Throughout this paper, we work with uniformly locally finite metric spaces (see Definition 1.1), also referred to as bounded geometry metric spaces. A classical example is provided by a Cayley graph of a finitely generated group with the shortest path metric. The coarse category $\mathbf{Coarse}$ is the category whose objects are uniformly locally finite metric spaces and whose morphisms are closeness classes of uniformly finite-to-one [^1] A map $f\colon X\to Y$ is uniformly finite-to-one if there exists $N\geq 1$ such that the preimage of every point of $Y$ consists of at most $N$ elements (see Section 1.1). Coarse equivalences are automatically uniformly finite-to-one. coarse maps. Isomorphisms in $\mathbf{Coarse}$ are called coarse equivalences. In the case of finitely generated groups equipped with word metrics, the notion of coarse equivalence coincides with quasi-isometry. We say that two uniformly locally finite metric spaces $(X,d)$ and $(Y,\partial)$ are (bijectively) coarsely equivalent if there exists a (bijective) coarse equivalence between them.
Coarse geometry is intrinsically connected to the theory of étale groupoids [^2] We refer the reader to [22] for a detailed introduction to étale groupoids and their $C^{*}$-algebras. . An étale groupoid is a topological groupoid $\mathcal{G}$ with unit space $\mathcal{G}^{(0)}$ such that the source and range maps $s,r\colon\mathcal{G}\to\mathcal{G}^{(0)}$ are local homeomorphisms. An étale groupoid $\mathcal{G}$ is said to be ample if its unit space $\mathcal{G}^{(0)}$ is totally disconnected. The theory of étale groupoids has attracted significant attention due to its deep connections with topological dynamics, number theory, and, in particular, operator algebras. To any étale groupoid $\mathcal{G}$ one can associate a $C^{*}$-algebra $C^{*}_{r}(\mathcal{G})$, called the reduced groupoid $C^{*}$-algebra of $\mathcal{G}$. In [14], Li showed that every classifiable [^3] By a “classifiable $C^{*}$-algebra” we mean a unital, separable $C^{*}$-algebra with finite nuclear dimension that satisfies UCT. simple $C^{*}$-algebra arises as the reduced groupoid $C^{*}$-algebra of a topologically principal (possibly twisted) étale groupoid. In [8], Crainic and Moerdijk introduced a homology theory for étale groupoids, which is now a substantial tool for analysing the $K$-theory of the reduced groupoid $C^{*}$-algebras arising from ample groupoids. Precisely, in [17] Matui showed that, for groupoids arising from shifts of finite type, the $K_{0}$-groups (respectively, $K_{1}$-groups) of the associated reduced groupoid $C^{*}$-algebras are isomorphic to the direct sum of the even (respectively, odd) homology groups. He conjectured that this phenomenon extends to more general ample groupoids.
Conjecture (HK conjecture).
$K_{*}\bigl(C^{*}_{r}(\mathcal{G})\bigr)\cong\bigoplus_{i=0}^{\infty}H_{2i+*}(\mathcal{G};\mathbb{Z}),$
where $H_{*}(\mathcal{G};\mathbb{Z})$ are the Crainic and Moerdijk homology groups of $\mathcal{G}$.
The HK conjecture has been verified for many classes of ample groupoids, in some cases even beyond the stated hypotheses. However, counterexamples to the HK conjecture were constructed by Scarparo [21] and Deeley [9]. It is well known that for any ample groupoid $\mathcal{G}$ and $i=0,1$ there exists canonical group homomorphism
called the $i$th comparison map (see Definition 1.15 for the definition of $\alpha_{0}$ in the case of coarse groupoids). At present, only partial results are known concerning the existence and properties of comparison maps in higher degrees.
For a uniformly locally finite metric space $(X,d)$, the authors of [23] constructed a Hausdorff, principal, ample groupoid $\mathcal{G}(X,d)$, called the coarse groupoid of $(X,d)$, whose unit space is the Stone-Cech compactification of $X$. The construction of the coarse groupoid yields a functor from the coarse category to the category of étale groupoids with generalised groupoid homomorphisms. The authors of [5] proved that the HK-conjecture holds for coarse groupoids associated with uniformly locally finite metric spaces of sufficiently small asymptotic dimension. Moreover, they showed that the groupoid homology of the coarse groupoid $\mathcal{G}(X,d)$ is canonically isomorphic to the uniformly finite homology (see Definition 1.10) of $(X,d)$, which will be denoted by $H_{i}^{\mathrm{uf}}(X;\mathbb{Z})$. The reduced groupoid $C^{*}$-algebra of a coarse groupoid associated to a uniformly locally finite metric space $(X,d)$ is more commonly known as the uniform Roe algebra $C^{*}_{u}(X)$. The latter has a simple description in terms of bounded operators on the Hilbert space $\ell^{2}(X)$. A bounded operator $T\in\mathcal{B}(\ell^{2}(X))$ is said to have controlled propagation if the quantity
@@
is finite. The uniform Roe algebra $C^{*}_{u}(X)$ is defined as the norm closure in $\mathcal{B}(\ell^{2}(X))$ of the *-algebra of bounded operators with controlled propagation. If $G$ is a finitely generated discrete group equipped with a word metric, then the uniform Roe algebra of $G$ is canonically isomorphic to the reduced crossed product $\ell^{\infty}(G)\rtimes_{r}G$. We refer to [20, Proposition 10.29] for an explicit construction of the isomorphism between $C^{*}_{u}(X)$ and $C^{*}_{r}(\mathcal{G}(X,d))$.
Uniform Roe algebras encode a substantial amount of coarse-geometric information about the underlying space. For example, a uniformly locally finite metric space $(X,d)$ has property A if and only if $C^{*}_{u}(X)$ is a nuclear $C^{*}$-algebra [23], and $(X,d)$ is amenable (as a metric space) if and only if $C^{*}_{u}(X)$ admits a tracial state [1]. This naturally leads to the question of whether the uniform Roe algebra fully determines the coarse geometry of the underlying space. The authors of [2] showed that the Morita equivalence class of the uniform Roe algebra is a complete invariant of coarse equivalence.
Theorem 0.1 (Baudier, Braga, Farah, Khukhro, Vignati, Willett, 2021).
@@
Let $(X,d)$ and $(Y,\partial)$ be uniformly locally finite metric spaces. The $C^{*}$-algebras $C^{*}_{u}(X)$ and $C^{*}_{u}(Y)$ are Morita equivalent if and only if $(X,d)$ and $(Y,\partial)$ are coarsely equivalent.
An analogous statement for isomorphisms of uniform Roe algebras and bijective coarse equivalences, known as “the bijective rigidity problem”, was a long-standing conjecture in coarse geometry. It was recently resolved by Vignati in [27].
Theorem 0.2 (Vignati, 2026).
@@
Let $(X,d)$ and $(Y,\partial)$ be uniformly locally finite metric spaces. The $C^{*}$-algebras $C^{*}_{u}(X)$ and $C^{*}_{u}(Y)$ are isomorphic if and only if $(X,d)$ and $(Y,\partial)$ are bijectively coarsely equivalent.
The $K$-theory of the uniform Roe algebra is of particular importance both from a mathematical perspective and for applications in physics. In [25], Špakula introduced the uniform coarse Baum–Connes conjecture, which asks whether a certain assembly map from the uniform $K$-homology of a metric space to the $K$-theory of the associated uniform Roe algebra is an isomorphism. On the other hand, the Baum–Connes conjecture for ample groupoids yields another assembly map, from the topological $K$-theory of the coarse groupoid to the $K$-theory of the associated uniform Roe algebra, which was conjectured [^4] Counterexamples to surjectivity were constructed in [12]. to be an isomorphism. It was shown in [11] that, when the uniformly locally finite metric space arises from a finitely generated group, the two conjectures are equivalent. In [13], Kubota used uniform Roe algebras as models for controlled topological phases of bulk and edge quantum systems. In particular, the $K$-theory of the uniform Roe algebra plays a decisive role in the study of the so-called bulk index.
@@
In this paper, we establish a connection between the bijective rigidity problem and the injectivity of the 0th comparison map appearing in the HK conjecture for coarse groupoids.
We say that a collection $\mathcal{C}$ of metric spaces is admissible (see Definition 3.1) if it is stable under taking subspaces and doublings [^5] On the level of coarse groupoids, a doubling corresponds to taking the product with $\mathcal{R}_{n}$, where $\mathcal{R}_{n}$ denotes the discrete groupoid associated to the full equivalence relation on $n$ points. See Section 1.1 for the coarse geometric definition. . For a property $\mathrm{P}$ of metric spaces, we say that an admissible collection $\mathcal{C}$ satisfies $\mathrm{P}$ if every element of $\mathcal{C}$ satisfies $\mathrm{P}$. For example, since property A passes to subspaces and doublings, the collection of metric spaces satisfying property A forms an admissible collection. The collection of all uniformly locally finite metric spaces, clearly, forms an admissible collection. The following lemma constitutes the main result of this paper.
Lemma A.
@@
    For every uniformly locally finite metric space $(X,d)$ in $\mathcal{C}$, the 0th comparison map $\alpha_{0}\colon H^{\mathrm{uf}}_{0}(X;\mathbb{Z})\to K_{0}\bigl(C^{*}_{u}(X)\bigr)$ is injective;
@@
    For any uniformly locally finite metric spaces $(X,d)$ and $(Y,\partial)$ in $\mathcal{C}$ such that $C^{*}_{u}(Y)\cong C^{*}_{u}(X)$, the coarse equivalence given by Theorem 0.1 is close to a bijective coarse equivalence.
By Theorem 0.2, the collection of all uniformly locally finite metric spaces satisfies the second condition of Lemma A; therefore, the first assertion holds unconditionally. Nevertheless, we formulate Lemma A as establishing a relationship between the two conjectural statements, as this viewpoint clarifies the techniques underlying our arguments. Moreover, we expect that the methods developed here can be adapted to broader classes of ample groupoids, potentially yielding new results on the injectivity of comparison maps and on rigidity phenomena for ample groupoids.
Theorem B.
The above theorem partially extends the results of [5], where the authors proved that $\alpha_{0}$ is split injective for uniformly locally finite metric spaces of small asymptotic dimension. To the best of our knowledge, no further results concerning the (split) injectivity of $\alpha_{0}$ for uniform Roe algebras are currently available. There exist many examples of uniformly locally finite metric spaces with infinite asymptotic dimension that are not coarsely connected, such as expander graphs. By contrast, any finitely generated group equipped with a word-length metric is coarsely connected. In this setting, for a finitely generated group $G$, the uniformly finite homology of $G$ is canonically isomorphic to the group homology of $G$ with coefficients in $\ell^{\infty}(G,\mathbb{R})$ (see [10, Proposition A.10]).
@@
Recent results of [19] show that the rational HK conjecture holds for all second-countable, locally compact, Hausdorff, ample groupoids that satisfy the rational Baum–Connes conjecture, and have torsion-free stabilisers. Coarse groupoids fall outside this framework: they are not second-countable and do not in general satisfy the Baum–Connes conjecture (see [12]). Nevertheless, by tensoring with $\mathbb{Q}$ in Theorem B, we immediately obtain unconditional split injectivity of the rational comparison map (see Remark 3.6).
@@
In Section 1, we introduce the necessary background from coarse geometry, uniform Roe algebras, and uniformly finite homology. We briefly recall the construction of the coarse groupoid from [23], sketch the main ideas underlying the proof of Theorem 0.1, review the results on uniformly finite homology, and construct the 0th comparison map.
Section 2 develops a framework of covering isometries for uniform Roe algebras, analogous to the construction for Roe algebras (see [30, Section 4.3]). Let $H$ be a separable Hilbert space. A uniform cover of a uniformly finite-to-one coarse map $f\colon X\to Y$ is an isometry
@@
In Section 3, we prove Lemma A. The key step in the proof of the implication $(1)\Rightarrow(2)$ is to show that, given an isomorphism of uniform Roe algebras $\Phi\colon C^{*}_{u}(X)\longrightarrow C^{*}_{u}(Y)$, the induced map in $K$-theory coincides with the map arising from the coarse equivalence $f_{\Phi}\colon X\longrightarrow Y$ provided by Theorem 0.1. Since $K_{0}(\Phi)$ preserves the class of the unit, it follows that
@@
modulo $\ker(\alpha_{0}^{Y})$. As the comparison maps are assumed to be injective, the result of [29, Theorem A] (see Theorem 1.12) completes the proof of the implication $(1)\Rightarrow(2)$. For the converse implication, to each positive function $h\in\ell^{\infty}(X,\mathbb{Z})$ we associate a subspace $X(h)$ of a doubling of $X$. We show that if $h_{1},h_{2}$ are positive functions such that
@@
This completes the proof of Lemma A. In the final section, we prove Theorem B and discuss further consequences of Lemma A.
Acknowledgements
The author is grateful to his advisors, Alessandro Vignati and Romain Tessera, for their advice on the presentation of this work. He also thanks Ján Špakula and the University of Southampton for their hospitality during his research visit. The author is indebted to Rufus Willett for valuable comments on Remark 4.3. This work was partially supported by the travel grant ECOST Actions CaLISTA (E-COST-GRANT-CA21109-15f1764d) and by the ANR grant ROAR (ANR-25-CE40-5029) .